A constant force of 245 N is applied at a 48.2° angle to a mass of 62.1 kg as shown below. If the mass moves at a constant speed of 3.28 ms–1, determine its coefficient of dynamic friction.  


1
Expert's answer
2022-04-13T09:20:42-0400

"\mu=\frac F{mg\sin\alpha}=0.9."